# Binary Autoencoder for Mechanistic Interpretability of Large Language Models

### v1

@@ 5.3 BAE is an Effective Interpretable Feature Extractor
As mentioned before, BAE effectively mitigates the two issues in feature extraction of typical SAE: dense features on inherently high-magnitude channels, which lower consistency among instances with the same feature activated (i.e., low interpretability scores in Step 3 of ComSem), and dead features on inherently low-magnitude channels, leaving many channels idle. In this section, we empirically confirm the mitigation of both issues by BAE with ComSem on BookCorpus (Zhu et al., 2015) against SAE variants trained on Pile (detailed in Appendix A.4) as shown in Table 1, where compared to all the baselines, BAE can extract the largest amount of active features from the LLM’s hidden states, with a considerable interpretability score. However, one can doubt that the interpretability score of BAE is “not SotA” in the Table 1. Our explanation is: BAE has a strong feature extracting ability, so that it captures some complex or vague features (refer to Appendix C) which are difficult to interpret by natural language from LLMs, which reduces the averaged interpretability scores. As shown in Appendix B, if some of the low-score features are removed, then BAE reaches the SotA of interpretability score and also remains the largest feature amounts.
Moreover, we conduct normalization [^7] I.e., linearly scale every channel value to a distribution over $h_{1}$ instances with mean value 0 and standard deviation $1$, to mitigate the influence of inherent activation magnitude of channels. to the $h_{1}$ of ReLU SAE for a rescaled activation magnitude, and evaluate as “ReLU SAE Resc.” in Table 1, where comparing to the vanilla ones, the interpretability scores are observably improved by the normalization rescaling. Since the training of SAE ignores cross-sample sparsity and its $h_{1}$ distribution is hard to estimate, the effectiveness of rescaling remains limited.
@@
To examine BAE’s effective suppression towards dense feature, we visualize the activation frequency distribution of each channel on the trained BAEs and SAEs as Fig. 6. In the visualizations, BAE features are sparsely activated with a left-leaning distribution, while some of the feature channels in the typical SAE keep high activation frequencies with a long-tail distribution, suggesting dense activations (Stolfo et al., 2025; Rajamanoharan et al., 2024b). Such results support our hypothesis that the minibatch-oriented entropy objective can mitigate the global density among input instances. Visualizations on more settings are provided in Appendix D.1.
@@ Conclusions.
As detailed in Appendix A.5, one can utilize BAE to compress vectors by storing only channel indices of $h_{1}$ where the burstiness against prior $\bar{h_{1}}$ exceeds a threshold. Such compression has an expected coding length for each instance [^8] Here, $\mathrm{L}_{\texttt{int}}$ or $\mathrm{L}_{\texttt{float}}$ denotes the number of bits required to encode a single index for the compressed coding or one float variable for an original feature element, and $\mathrm{H}[h_{0}]$ is the entropy estimated by the BAE from the distribution of $h_{0}$, $d$ is the dimensionaility of $h_{0}$. The $\mathrm{L}_{\texttt{int}}$ is usually less than $\mathrm{L}_{\texttt{float}}$, and $\mathrm{H}[h_{0}]$ is also usually less than $d$ (as demonstrated in Fig. 4, with $d=2048$ and $\mathrm{H}[h_{0}]<700$). $\mathrm{L}_{\texttt{int}}\mathrm{H}[h_{0}]$, outdistances the original $\mathrm{L}_{\texttt{float}}d$. Reconstruction is done by flipping $\mathrm{round}(\bar{h_{1}})$ at the stored indices and projecting through $W_{\text{out}}$. We evaluate this compression with a threshold $\log 0.5$ (refer Appendix A.5 for details), as shown in Table 2, achieving compression rates as low as 1% with low MSE, indicating both effective compression and reconstruction quality. However, the reconstruction remains distorted in the radial direction, manifesting as low cosine similarity with the originals, likely arising from the $L_{2}$ loss of BAE, which ignores radial information. Since this method constitutes lossy compression and falls outside the paper’s main focus, we do not present it as a core contribution but a potential application. For reliable and efficient storage or transmission, specialized objectives beyond simple regression are needed to preserve essential information better.
@@
Since BAE represents feature activations with binary values rather than continuous ones, inherently continuous attributes (such as RGB values (Engels et al., 2025; Modell et al., 2025)) may be discretized into multiple features, which is reasonable since these features cannot be represented with 1-bit, so decomposing them into multiple features facilitates the correct estimation of the entropy. On the other hand, due to the narrow radial distribution of Transformer hidden states (Gao et al., 2019; Ethayarajh, 2019), different features may share the same direction but differ in distance, which is difficult to distinguish in continuous SAEs, whereas BAE avoids this issue by discretizing them into atomic units. Also, refer to Table 1, compared to baselines, the number of discretized features extracted by BAE significantly increase against layers, with SAE features remain constant, suggesting that continuous features are more likely to emerge in later layers, which are split into more discrete features by BAE, consistent with previous works (Jawahar et al., 2019; Chen et al., 2023; Allen-Zhu & Li, 2023; Liu et al., 2024b), still needs further exploration.
@@
(1) Non-zero Reconstruction Loss. In our experiments, the regression loss remains notable, which may cause a performance degradation. Future work can be devoted to modifying the structure of BAE, especially by more advanced binarization functions (Wang et al., 2018; Vargas et al., 2024) and gradient estimation methods (Darabi et al., 2019; Yang et al., 2019); also, more scaled training sets for BAE can be utilized in an industrial scenario. (2) Natural Language Feature Interpretation Based on Tokens. In this paper, ComSem, and also the traditional automatic feature interpretation methods, are all based on the natural language interpretation of tokens where the features originate. However, the contextualization induced by Transformer layers may cause the semantics of the hidden states to deviate from the original token, but the method family overlooks such an effect. Future works can be devoted to directly decoding the semantics from the extracted atomized features by tools such as LogitLens (Joseph Bloom, 2024) or PatchScopes (Ghandeharioun et al., 2024). Also, not all the features can be interpreted by natural languages, such as Task Vectors (Hendel et al., 2023; Kharlapenko et al., 2024). We discuss this point in Appendix C. (3) Investigation Scope. Due to computational limits, we evaluate BAE only on Llama 3.2-1B and 3B, leaving large-scale tests for future work. Nevertheless, we believe this paper provides a sufficient prototype of BAE for addressing dense/dead feature issues in SAE.
@@ Acknowledgments
The Use of Large Language Models
In this paper, LLMs are used to polish writing and also help the experiment as described in §5 under the LLM-as-a-judge framework (Gu et al., 2024).
Appendices
@@ Appendix A Experiment Details
We generate all $65536$ samples, with $n=52428$ training samples and the remaining as validation samples, with non-mentioned hyperparameters kept as the default. After the training, we run BAE again on all $65536$ samples before for binary encoding $h_{1}$, and calculate the marginal entropy on these $h_{1}$ as mentioned in Eq. 4.
@@
In the experiment of §4.2, we generate $n=209715$ data samples for the training of BAE, with more $52429$ data samples as the validation set from the Pile-train split, with non-mentioned hyperparameters kept as the default. Especially, we filter out all the input sentences to the Llama 3.2-1B with a token length less than 1024 to keep the length of hidden state sets among all positions aligned. After the training, we run BAE again on all $262144$ samples before for binary encoding $h_{1}$, and calculate the marginal entropy on these $h_{1}$ as mentioned in Eq. 4.
@@
ICL (Radford et al., 2019; Dong et al., 2022) typically utilizes concatenations of input($x$)-answer($y$) pairs (with amount $k$, called demonstrations) to define a task, and requires the LM to generate the answer of the last input ($x_{q}$, called query) similar to the demonstrations. The inputs of ICL are built like $x_{1},y_{1},x_{2},y_{2},\dots,x_{k},y_{k},x_{q}]$. In practice, the input samples used in the experiments are built on the toolkit StaICC (Cho & Inoue, 2025), similar to the instance below:
sentence: a genuinely moving and wisely unsentimental drama . sentiment: positive sentence: laughs – sometimes a chuckle , sometimes a guffaw and , to my great pleasure , the occasional belly laugh sentiment: positive sentence: that the entire exercise has no real point sentiment: negative sentence: 90 punitive minutes of eardrum-dicing gunplay , screeching-metal smashups , and flaccid odd-couple sniping sentiment: negative sentence: amid the new populist comedies that underscore the importance of family tradition and familial community sentiment: positive sentence: freakshow sentiment: negative sentence: a taste for the quirky sentiment: positive sentence: rustic , realistic , and altogether creepy sentiment: positive sentence: assured direction and complete lack of modern day irony sentiment:
@@
We sample $262144$ ICL input samples with $k\in\{0,1,2,4,8\}$, and extract the hidden states of the last token (i.e., “:”) from layer $\{0,2,4,6,8,10,12,14,16\}$, with $209715$ samples for the training, and the remaining for validation. After the training, we run BAE again on all $262144$ samples before for binary encoding $h_{1}$, and calculate the marginal entropy on these $h_{1}$ as mentioned in Eq. 4.
@@
  • Top-$k$ SAE (Bussmann et al., 2024). The feedforward calculation of Top-$k$ SAE is defined as:
@@
  • Gated ReLU SAE (Rajamanoharan et al., 2024a). The feedforward calculation of Gated ReLU SAE is defined as:
@@
  • Transcoder (Dunefsky et al., 2024). The feedforward calculation of Transcoder is defined as:
@@
The pseudocode of ComSem is shown in Algorithm 1, with default parameters $N=8192$, $n_{I}=5$, $n_{T}=8$, $k=10$. The evaluations are conducted on the train split of BookCorpus, which is different from the training set of BAE/SAEs (Pile). The $\bar{h_{1}}$ is averaged among all the input instances of the evaluation set. The instructions for calling the backend LLMs are shown in Fig. 20 and 21, with queries’ format the same as the given examples. We utilize GPT-4.1 and GPT-4.1-mini (Achiam et al., 2023) as the backend LM.
@@ A.5 Vector Compression by BAE
    Given the input $h_{0}$ to be compressed, we first compute its binary encoding $h_{1}$ by applying the input projection $W_{\text{in}}$ and binarize function $\Gamma$ of a BAE, and then measure the burstiness $\beta$ with respect to the prior distribution $\bar{h_{1}}$, as described in §5.2 and Fig. 10.
@@
    Given a threshold [^10] Notice that the threshold greater than 0 is trivial, given the $h_{1}\in\{0,1\}^{d^{\prime}}$, $\bar{h}_{1}\in[0,1]$, so that all the elements of $\log|h_{1}-\bar{h}_{1}|\leqslant 0$. And burstiness less than $\log 0.5$ causes wrong flip of bits in Step 3. $B\in[\log 0.5,0]$, we bookkeep only the channel indices $i$ for which $\beta_{i}>B$, using these indices as the compressed representation of $h_{0}$.
@@
    To reconstruct $h_{0}$ from the bookkept indices, we round $\bar{h_{1}}$ to a binary vector, flip the bits of $\mathrm{round}(\bar{h_{1}})$ at the stored indices, and then pass the flipped binary vector through $W_{\text{out}}$ (also the bias term) to obtain the recovered $h_{0}$.
@@
We sample $117864$ hidden state vectors from the specific layers on BookCorpus, then utilize the BAE trained on Pile for the compression, following the processing above. The $\bar{h_{1}}$ is averaged among all the input instances of the compressed set.
Appendix B Interpretability Score Distribution: Selecting a Better Interpretable Feature Set
We visualize the ComSem interpretability score distribution among features of BAE and baselines in Fig. 2226, where compared to all the baselines, BAE significantly increased the amount of both high-score and low-score features. As shown in Appendix C, these low-score features are non-literal, that is, the semantics of these features can not be interpreted by the token inputs or outputs in ComSem, so that, despite serving as an effective proof of BAE’s stronger extraction capability, harm the average interpretability score among all the activated features.

### v2

@@ 5.3 BAE is an Effective Interpretable Feature Extractor
As mentioned before, BAE effectively mitigates the two issues in feature extraction of typical SAE: dense features on inherently high-magnitude channels, which lower consistency among instances with the same feature activated (i.e., low interpretability scores in Step 3 of ComSem), and dead features on inherently low-magnitude channels, leaving many channels idle. In this section, we empirically confirm the mitigation of both issues by BAE with ComSem on 2 backend evaluation LLMs and BookCorpus (zhu2015aligning) against SAE variants trained on Pile (detailed in Appendix A.4) as shown in Table 1, where compared to all the baselines, BAE can extract the largest amount of active features from the LLM’s hidden states, with a considerable interpretability score. However, one can doubt that the interpretability score of BAE is “not SotA” in the Table 1. Our explanation is: BAE has a strong feature extracting ability, so that it captures some complex features (refer to Appendix F) which are difficult to interpret by natural language from LLMs, reducing the interpretability scores. As shown in Appendix C, if some of the low-score features are removed, then BAE reaches the SotA of interpretability score and also remains the largest feature amounts. Moreover, we conduct normalization [^5] I.e., linearly scale every channel value to a distribution over $h_{B}$ instances with mean value 0 and standard deviation $1$, to mitigate the influence of inherent activation magnitude of channels. to the $h_{B}$ of ReLU SAE for a rescaled activation magnitude, and evaluate as “ReLU SAE Resc.” in Table 1, where comparing to the vanilla ones, the interpretability scores are observably improved by the normalization rescaling. Since SAE training ignores cross-sample sparsity and its $h_{B}$ distribution is hard to estimate, the effectiveness of rescaling remains limited.
Figure 5: Feature activation frequency distribution of Layer 11 (more layers in Appendix G.3).
Information Loss by Binarization.
One suspicion is that the $\Gamma$ causes information loss from the $h_{0}$, which may harm the feature extraction. To evaluate such information loss, we calculate the self-regression loss ($\mathcal{L}_{r}$) on the validation set, as shown in Table 1, where no significant difference of BAE compared to baselines can be observed. This indicates that BAE at least does not incur more severe information loss than the baselines. Furthermore, we attempt to increase the decoding bits to reduce such information loss in §6.
@@
To examine BAE’s effective suppression towards dense feature, we visualize the activation frequency distribution of each channel on the trained BAEs and SAEs as Fig. 5. In the visualizations, BAE features are sparsely activated with a left-leaning distribution, while some of the feature channels in the typical SAE keep high activation frequencies with a long-tail distribution, suggesting dense activations (stolfo2025antipodal; rajamanoharan2024jumping). Such results support our hypothesis that the minibatch-oriented entropy objective can mitigate the global density among input instances. Visualizations on more settings are provided in Appendix G.1.
Feature Diversity.
As an intuitive measurement of feature diversity, we directly compute the spectral norm of the covariance matrix over the row vectors in $W_{\text{out}}$, with the results reported in Table 1. Among these results, BAE exhibits a markedly larger covariance, confirming its ability to extract diverse and non-redundant features.
5.4 Case Analysis: BAE Extract Better Steering Vectors
As empirical evidence that BAE can extract effective features, we reconstruct the ICL task vector (hendel-etal-2023-context) using the features extracted by BAE and SAE baselines (sae_taskvectors). In detail, following the framework proposed by cho-hockenmaier-2025-toward, we construct an 8-shot sample set as the “good samples”, and a 0-shot sample set as the “bad samples” from MR (MR), both equipped with 512 samples. We input both sets respectively into LMs, getting the hidden states of the last token from Layer 11 of Llama 3.2-1B, and encoding the hidden states by BAE or SAE baselines. Then, we calculate $h_{D}=\bar{h}_{B}^{+}-\bar{h}_{B}^{-}$, where the $\bar{h}_{B}^{+}$ and $\bar{h}_{B}^{-}$ are the mean hidden activation in these two sets produced by BAE or SAE baselines, and only keep the elements at the top 256 in absolute value (i.e., zeroing the non-significant values) as $\tilde{h_{D}}$. $\gamma\tilde{h_{D}}W_{\text{out}}$ is calculated as the steering vector with $\gamma$ as injection strength, and injected into the original Layer 11 of the last-token hidden states from the 0-shot test samples, with accuracy shown in Fig. 6. The results show that the BAE-mediated steering vector most effectively improves zero-shot accuracy, serving as a case study demonstrating that BAE can extract more effective features from LLMs.
Figure 6: Steering vector reconstruction results on MR, Llama 3.2-1B Layer 11.
6 More-bit Decoding
As mentioned before, although the self-regression loss induced by BAE is not significant compared to the baselines, it can be beneficial to reduce such loss. Therefore, based on the framework of BAE, in this section, as shown in Fig. 1, we propose a plug-in module to aggregate multiple bits of $h_{B}$ into real variables before decoding by the $W_{\text{out}}$. In detail, for each $C$ binary elements in $h_{B}\in\{0,1\}^{d^{\prime}}$ (i.e., from the $iC$-th elements to the [(i+1)C1][(i+1)C-1]-th elements of $h_{B}$), we transfer the binary bits into a real vector as:
$h_{R}=\frac{\sum_{j=0}^{C-1}2^{j}h_{B,iC+j}}{{2^{C}-1}},$ (7)
then, we decode $h_{R}$ by $W_{\text{out}}\in\mathbb{R}^{d^{\prime}/C\times d}$ for the output $\hat{h_{0}}$. Such a method effectively improves the numerical precision of BAE, and is therefore expected to mitigate the information loss while still retaining a binary encoding to support entropy computation. The entropy loss is calculated on the binary encoding $h_{B}$, whereas feature interpretation is conducted on the real-valued $h_{R}$.
Evaluation results with $d^{\prime}/C=4d$ are shown in Table 2, where more-bit decoding consistently reduces self-regression loss while decreasing the number of activation features, yet remains superior to the SAE baseline. As we will discuss in §7, BAE decomposes intrinsically continuous features into multiple atomic ones, whereas more-bit decoding merges them into a single feature, which leads to a reduction in the total number of features. Such a result demonstrates that more-bit decoding serves as an effective trade-off handle between feature numbers and information loss. Moreover, to reduce the information loss of BAE, future work can be devoted to modifying the structure of BAE, especially by more advanced binarization functions (wang2018two; vargas2024biper) and gradient estimation methods (darabi2019bnn; yang2019quantization).
Table 2: More-bit decoding results on Llama 3.2-1B Layer 11. Notation similar to Table 1. Evaluated on GPT 4.1-mini.
Bits ($C$) $\mathcal{L}_{r}$ Act. FI# Score
[…]
@@ Conclusions.
As detailed in Appendix A.5, one can utilize BAE to compress $h_{0}$ set by storing only channel indices of $h_{B}$ where the burstiness against prior $\bar{h_{B}}$ exceeds a threshold. Such compression has an expected coding length for each instance [^6] Here, $\mathrm{L}_{\texttt{int}}$ or $\mathrm{L}_{\texttt{float}}$ denotes the number of bits required to encode a single index for the compressed coding or one float variable for an original feature element, and $\mathrm{H}[h_{0}]$ is the entropy estimated by the BAE from the distribution of $h_{0}$, $d$ is the dimensionaility of $h_{0}$. The $\mathrm{L}_{\texttt{int}}$ is usually less than $\mathrm{L}_{\texttt{float}}$, and $\mathrm{H}[h_{0}]$ is also usually less than $d$ (as demonstrated in Fig. 3, with $d=2048$ and $\mathrm{H}[h_{0}]<700$). $\mathrm{L}_{\texttt{int}}\mathrm{H}[h_{0}]$, outdistances the original $\mathrm{L}_{\texttt{float}}d$. Reconstruction is done by flipping $\mathrm{round}(\bar{h_{B}})$ at the stored indices and projecting through $W_{\text{out}}$. We evaluate this compression with a threshold $\log 0.5$ (refer Appendix A.5), as shown in Table 3, achieving compression rates as low as 1% with low MSE, indicating both effective compression and reconstruction quality. However, the reconstruction remains distorted in the radial direction, manifesting as low cosine similarity with the originals, likely arising from the $L_{2}$ loss of BAE, which ignores radial information. Since this method constitutes lossy compression and falls outside the paper’s main focus, we do not present it as a core contribution but a potential application. For reliable and efficient storage or transmission, specialized objectives beyond simple regression are needed to preserve essential information better.
@@
Since BAE represents feature activations with binary values rather than continuous ones, inherently continuous attributes (such as RGB values (engels2025not; modell2025origins)) may be discretized into multiple features, which is reasonable since these features cannot be represented with 1-bit, so decomposing them into multiple features facilitates the correct estimation of the entropy. On the other hand, due to the narrow radial distribution of Transformer hidden states (gao2018representation; ethayarajh-2019-contextual), different features may share the same direction but differ in distance, which is difficult to distinguish in continuous SAEs, whereas BAE avoids this issue by discretizing them into atomic units. Also, refer to Table 1, compared to baselines, the number of discretized features extracted by BAE significantly increase against layers, with SAE features remain constant, suggesting that continuous features are more likely to emerge in later layers, which are split into more discrete features by BAE, consistent with previous works (jawahar-etal-2019-bert; chen2023which; allen2023physics; liu-etal-2024-fantastic), still needs further exploration.
@@
(1) Natural Language Feature Interpretation Based on Tokens. In this paper, ComSem, and also the traditional automatic feature interpretation methods, are all based on the natural language interpretation of tokens where the features originate. However, the contextualization induced by Transformer layers may cause the semantics of the hidden states to deviate from the original token, but the method family overlooks such an effect. Future works can be devoted to directly decoding the semantics from the extracted atomized features by tools such as LogitLens (bloom2024understandingfeatureslogitlens) or PatchScopes (pmlr-v235-ghandeharioun24a). Also, not all the features can be interpreted by natural languages, such as Task Vectors (hendel-etal-2023-context; sae_taskvectors). We discuss this point in Appendix F. (2) Investigation Scope. Due to computational limits, we evaluate BAE only on Llama 3.2-1B and 3B, leaving large-scale tests for future work. Nevertheless, we believe this paper provides a sufficient prototype of BAE for addressing dense/dead feature issues in SAE.
@@ Acknowledgments
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Experiment Details
We generate all $65536$ samples, with $n=52428$ training samples and the remaining as validation samples, with non-mentioned hyperparameters kept as the default. After the training, we run BAE again on all $65536$ samples before for binary encoding $h_{B}$, and calculate the marginal entropy on these $h_{B}$ as mentioned in Eq. 4.
@@
In the experiment of §4.2, we generate $n=209715$ data samples for the training of BAE, with more $52429$ data samples as the validation set from the Pile-train split, with non-mentioned hyperparameters kept as the default. Especially, we filter out all the input sentences to the Llama 3.2-1B with a token length less than 1024 to keep the length of hidden state sets among all positions aligned. After the training, we run BAE again on all $262144$ samples before for binary encoding $h_{B}$, and calculate the marginal entropy on these $h_{B}$ as mentioned in Eq. 4.
@@
ICL (radford2019language; dong2022survey) typically utilizes concatenations of input($x$)-answer($y$) pairs (with amount $k$, called demonstrations) to define a task, and requires the LM to generate the answer of the last input ($x_{q}$, called query) similar to the demonstrations. The inputs of ICL are built like $x_{1},y_{1},x_{2},y_{2},\dots,x_{k},y_{k},x_{q}]$. In practice, the input samples used in the experiments are built on the toolkit StaICC (cho2025staicc), similar to the instance below:
sentence: a genuinely moving and wisely unsentimental drama . sentiment: positive
sentence: laughs – sometimes a chuckle , sometimes a guffaw and , to my great pleasure , the occasional belly laugh sentiment: positive
sentence: that the entire exercise has no real point sentiment: negative
sentence: 90 punitive minutes of eardrum-dicing gunplay , screeching-metal smashups , and flaccid odd-couple sniping sentiment: negative
sentence: amid the new populist comedies that underscore the importance of family tradition and familial community sentiment: positive
sentence: freakshow sentiment: negative
sentence: a taste for the quirky sentiment: positive
sentence: rustic , realistic , and altogether creepy sentiment: positive
sentence: assured direction and complete lack of modern day irony sentiment:
@@
We sample $262144$ ICL input samples with $k\in\{0,1,2,4,8\}$, and extract the hidden states of the last token (i.e., “:”) from layer $\{0,2,4,6,8,10,12,14,16\}$, with $209715$ samples for the training, and the remaining for validation. After the training, we run BAE again on all $262144$ samples before for binary encoding $h_{B}$, and calculate the marginal entropy on these $h_{B}$ as mentioned in Eq. 4.
@@
  • Top-$k$ SAE (bussmann2024batchtopk). The feedforward calculation of Top-$k$ SAE is defined as:
@@
  • Gated ReLU SAE (rajamanoharan2024improving). The feedforward calculation of Gated ReLU SAE is defined as:
@@
  • Transcoder (dunefsky2024transcoders). The feedforward calculation of Transcoder is defined as:
@@
The pseudocode of ComSem is shown in Algorithm 1, with default parameters $N=8192$, $n_{I}=5$, $n_{T}=8$, $k=10$. The evaluations are conducted on the train split of BookCorpus, which is different from the training set of BAE/SAEs (Pile). The $\bar{h_{B}}$ is averaged among all the input instances of the evaluation set. The instructions for calling the backend LLMs are shown in Fig. 22 and 23, with queries’ format the same as the given examples. We utilize GPT-4.1 and GPT-4.1-mini (achiam2023gpt) as the backend LM.
@@ A.5 Vector Compression by BAE
    Given the input $h_{0}$ to be compressed, we first compute its binary encoding $h_{B}$ by applying the input projection $W_{\text{in}}$ and binarize function $\Gamma$ of a BAE, and then measure the burstiness $\beta$ with respect to the prior distribution $\bar{h_{B}}$, as described in §5.2 and Fig. 10.
@@
    Given a threshold [^8] Notice that the threshold greater than 0 is trivial, given the $h_{B}\in\{0,1\}^{d^{\prime}}$, $\bar{h}_{1}\in[0,1]$, so that all the elements of $\log|h_{B}-\bar{h}_{1}|\leqslant 0$. And burstiness less than $\log 0.5$ causes wrong flip of bits in Step 3. $B\in[\log 0.5,0]$, we bookkeep only the channel indices $i$ for which $\beta_{i}>B$, using these indices as the compressed representation of $h_{0}$.
@@
    To reconstruct $h_{0}$ from the bookkept indices, we round $\bar{h_{B}}$ to a binary vector, flip the bits of $\mathrm{round}(\bar{h_{B}})$ at the stored indices, and then pass the flipped binary vector through $W_{\text{out}}$ (also the bias term) to obtain the recovered $h_{0}$.
@@
We sample $117864$ hidden state vectors from the specific layers on BookCorpus, then utilize the BAE trained on Pile for the compression, following the processing above. The $\bar{h_{B}}$ is averaged among all the input instances of the compressed set.
Appendix B Evaluating Entropy Estimation on LLM Hidden States Set
To extend our discussion in §4.1, in this section, we focus on evaluating the entropy evaluation on benchmarks extracted from real LLMs. In detail, similar to §4.1, given a ground-truth entropy $r$, we sample $n$ $r$-length token sequences, with elements randomly sampled from two given tokens (in our implementation, tokens with index 0 and 1). From such a sampling method, we can get an input sequence set with entropy $r$. Input these sequences into LMs (Llama 3.2-1B in our implementation) and extract the hidden states from the $l$-th layer (Layer 12 in our implementation), we can get a hidden states set with entropy not larger than rr [^9] Notice that in §4.2, we show that during the forward pass of an LM, the hidden state of a token may not contain all information from its preceding context; therefore, hidden states on a single position may carry reduced information.. Repeat the training-and-evaluation processing shown in §4.1, we calculate the entropy with varios $r$ as shown in Fig. 11, where the results are consistent with our inference. But this also indicates that such methods based on LM hidden states cannot serve as a strong benchmark and should only be considered a complement to our synthetic approach.
Appendix C Interpretability Score Distribution: Selecting a Better Interpretable Feature Set
Figure 11: Evaluation results on LM-extracted hidden states.
We visualize the ComSem interpretability score distribution among features of BAE and baselines in Fig. 2428, where compared to all the baselines, BAE significantly increased the amount of both high-score and low-score features. As shown in Appendix F, these low-score features are non-literal, that is, the semantics of these features can not be interpreted by the token inputs or outputs in ComSem, so that, despite serving as an effective proof of BAE’s stronger extraction capability, harm the average interpretability score among all the activated features.
Appendix D Natural Low-entropy Tendency of BAE Training